OpenStudy (anonymous):

prove that if n is a sum of two squares of non-negative integers then 2n is also a sum of 2 non-negative integers.

OpenStudy (anonymous):

its easy i think

OpenStudy (anonymous):

n=m^2+p^2 2n=2m^2+p^2=(m+p)^2+(p-m)^2

OpenStudy (anonymous):

if m>p then 2n=(m+P)^2+(m-p)^2

OpenStudy (anonymous):

so saubhik what was ur rank in jee?

OpenStudy (anonymous):

453 thnx for solving.. :)

OpenStudy (anonymous):

which year?

OpenStudy (anonymous):

this yr

OpenStudy (anonymous):

wht stream?

OpenStudy (anonymous):

engg

OpenStudy (anonymous):

arre yaar which stream did u choose>?

OpenStudy (anonymous):

counselling didnt start

OpenStudy (anonymous):

this rank is in IITjee or WBjee??

OpenStudy (anonymous):

i think wbjee

OpenStudy (anonymous):

yes I didnt get IIT

OpenStudy (anonymous):

so are u going to shibpur(BESU)

OpenStudy (anonymous):

would love to...but can I get anything in Jadavpur...i know my rank is poor...to get nything in Jadavpur...

OpenStudy (anonymous):

u can get com sc. mechanical. civil, may be elec

OpenStudy (anonymous):

which do u suggest is better ?

OpenStudy (anonymous):

whats ur choice?

OpenStudy (anonymous):

i think mech in jadavpur will be good? wat u think?

OpenStudy (anonymous):

u can get. my assumption is ..

OpenStudy (anonymous):

@saubhik i did ur another problem that says m^2<n<(m+1)^2 did u check it??